Problem:
Rose’s garden is in the shape of a trapezoid. If the height of the trapezoid is 16 m, one base is 20 m, and the area is 224 m2, find the length of the other base.
Area of a trapezoid= A=1/2h(B+b)
Hints:
B and b are 2 different variables. B does not = b
B and b are both bases of the trapezoid. (A trapezoid has 2 bases)
B is 20 m; you must find b.
You may want to use your answer to question 24 above to solve this problem, BUT IT IS NOT ACTUALLY NECESSARY; it just makes it slightly easier.
Be sure to check your answer. Put your answer back into the formula for the area of a trapezoid. Does the garden area still equal 224 m2?

Answer by ankor@dixie-net.com(22740)

Rose’s garden is in the shape of a trapezoid. If the height of the trapezoid is 16 m, one base is 20 m, and the area is 224 m2, find the length of the other base.
Area of a trapezoid= A=1/2h(B+b)
:
.5h(B+b = A
:
.5(16)(20+b) = 224; find b
:
8(20+b) = 224
160 + 8b = 224
8b = 224 - 160
8b = 64
b = 64/8
b = 8
:
Check:
.5(16)(20+8) =
8 * 28 = 224